2/3 DIVIDE 1 3/y = 11/21 find y
Answer by girl_from_moscow(7)

FIRST, are you kidding, or are you really asking this!!!
here is how you do this
1/3y=11/21
you probably know how to divide fractions but here is a short explanation.
if you want to dived a fraction by a fraction, then you need to switch the numerator with the denominator. then instead of dividing it you multiply your two fractions.
so 11/21 multiply by 3/1
which equals to 33/11 but you can always simplify the fraction by 3
so final answer= 11/7
ANSWER= 11/7 OR 1 and 4/7
HOPE YOU NEVER HAVE PROBLEMS WITH FRACTIONS AGAIN!! :DD